You own a bond that pays $110 in annual​ interest, with a ​$1,000 par value. It matures in 10 years. Your required rate of return is 12 percent.

a. Calculate the value of the bond.

b. How does the value change if your required rate of return​ (1) increases to 14 percent or​ (2) decreases to77 ​percent?

c. Explain the implications of your answers in part ​(b​) as they relate to interest rate​ risk, premium​ bonds, and discount bonds.

d. Assume that the bond matures in 33 years instead of 10 years. Recompute your answers in part (b​).

e. Explain the implications of your answers in part (d​) as they relate to interest rate​ risk, premium​ bonds, and discount bonds.
Answer #1

1.
=110/12%*(1-1/1.12^10)+1000/1.12^10=943.497769715891

2.
=110/14%*(1-1/1.14^10)+1000/1.14^10=843.516530611193

3.
=110/7%*(1-1/1.07^10)+1000/1.07^10=1280.9432616373

4.
Premium bonds have lower yield and thus have higher interest rate risk
Discount bonds have higher yield and thus have lower interest rate risk

5.
=110/14%*(1-1/1.14^33)+1000/1.14^33=788.553074787638

6.
=110/7%*(1-1/1.07^33)+1000/1.07^33=1510.15160067053

7.
Premium bonds have lower yield and thus have higher interest rate risk
Discount bonds have higher yield and thus have lower interest rate risk